OpenStudy (anonymous):

Can someone help me solve this please??? I need to solve for y!!! -5x+3y=-9

OpenStudy (anonymous):

Try to get the y term on one side of the equals sign, and have the x term, and the constant on the other side of the equals sign.

OpenStudy (anonymous):

\[y=-3+\left(\begin{matrix}5 \\ 3\end{matrix}\right)x\]

OpenStudy (anonymous):

you always want to isolate the variable you want to solve for, in this case its y. so you want to move the -5x to the other side turning it onto 3y= -9 +5x. the you want to divide each of those terms by three to isolate the y and you have your answer

OpenStudy (anonymous):

so is it y= -5/3x-3

OpenStudy (anonymous):

the 5/3 is positive

OpenStudy (anonymous):

because it was the 5x was negaztive on one side when you move it over it turns positive, then when you divide by three its just 5/3

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

no the answer was wrong I think the answer was y=3/5x-3

OpenStudy (anonymous):

-5x+3y=-9 ===> y= (5/3)x -3

OpenStudy (anonymous):

@desibaybee , your answer is incorrect.

OpenStudy (amistre64):

-5x+3y=-9 /3 /3 /3 ----------- -5/3x +y = -3 +5/3x +5/3x ---------------- y = (5/3)x -3